# WW-FL: Secure and Private Large-Scale Federated Learning

### v3

@@ 5.2. New Robust Aggregation Scheme
When computing the proposed trimmed mean variant, we are only interested in the indices of the outliers and therefore only perform the swap operations on a list of indices. This is done to minimize the compute operations and thereby reduce time and communication. After identifying which gradients were most often detected as outliers and then computing the set of benign indices, we finally compute the sum over those while preserving privacy. The procedure is shown in Lst. LABEL:listing:sumofbenign.
@@
3 counts = crypten.stack([indices for _ in range(num_bins)],dim=1).eq(indices_list).
4 sum(dim=0)
@@
6
7def exclude_topk_indices(self, indices, k):
@@
11
12benign_indices = self.exclude_topk_indices(
@@
15
16all_indices = crypten.cryptensor(torch.arange(0, n).repeat(benign_indices.size(0), 1)).t()
17agg_sum = gradients.matmul(all_indices.eq(benign_indices)).
18 sum(dim=1)
19
20return agg_sum / (n - self.exclude_topk)
@@
FLTrust (Cao et al., 2021) uses a root data set to calculate the server model. To be most effective, the server model must be similar to the benign models and therefore the root dataset must be representative for the whole dataset. Therefore, we sample (like for all clients) 200 data points for that dataset.
Table 8. Robust aggregation parameters.
Parameter FL WW-FL
@@
Figure 5. Validation accuracy for FL and WW-FL when training ResNet9 on CIFAR10 with FedAvg, FLTrust, and trimmed mean as aggregation schemes under DLF attack for three different poison rates (top: equally distributed, bottom: focused setting).
5.3. Evaluation
We now empirically evaluate the impact of data-poisoning attacks on WW-FL considering different (robust) aggregation schemes. For this, we implement all four attacks—RLF (Xiao et al., 2012), TLF (Tolpegin et al., 2020), SLF (Fang et al., 2020; Shejwalkar et al., 2022), DLF (Shejwalkar et al., 2022)—in our framework and add CrypTen-based implementations of three robust aggregation schemes to our prototype implementation. Using the setup described in §4, we answer the following questions next:
  • Q4:
@@ 5.3. Evaluation
Q4 – Attack Impact. We consider three poison rates (0.01, 0.1, 0.2) and distributions of attackers: in the equally-distributed setting, we assume that on expectation each cluster has the same number of malicious clients; in the focused setting, we assume that malicious clients are concentrated on as few clusters as possible while there is still an honest majority in each cluster (a standard assumption in FL); finally, in the unrealistic cluster-focused setting, we see what happens if we lift the honest-majority assumption and concentrate all malicious clients in as few clusters as possible. In Fig. 5, we study the effectiveness of the most powerful DLF (Shejwalkar et al., 2022) attack on both regular FL and WW-FL when training ResNet9 on CIFAR10 in the equally distributed and focused setting. Additional experimental outcomes, including evaluation over the cluster-focused setting, are given in §A.3.2.
@@
Q5 – Robust Aggregation in MPC. We evaluate the run-time and communication overhead of our FLTrust and TM implementation in CrypTen in Tab. 9. The run-time overhead for both robust aggregation schemes compared to FedAvg is four to five orders of magnitude higher (ranging from 5 to 30 minutes). Also, FLTrust requires 5$\times$ more run-time and communication than TM. Given that both produce fairly similar results when applied to WW-FL, the overhead for FLTrust seems not justified.
Table 9. Communication (Comm. in MB) and run-time (Time in seconds) for various aggregation schemes, including our Trimmed Mean (TM) variant with sample sizes 10, 100, and 1000 in CrypTen.
@@
Q6 – Trimmed Mean Variant. In Fig. 6, we compare the effectiveness of our TM variant to the original TM (Yin et al., 2018) for three sample sizes (10, 100, and 1000). It turns out that our heuristic approach barely reduces the effectiveness, even with aggressive parameters. In fact, in the focused setting, the TM variant outperforms the original. This is because our variant completely excludes updates of (poisoned) outliers, whereas in regular trimmed mean, those poisoned updates might still be considered for some coordinates. Results for all other settings are presented in §A.3.3.
Figure 6. Accuracy of trimmed mean (TM) and our variant (with sample sizes 10, 100, and 1000) against focused DLF attacks on WW-FL at 0.2 poison rate for ResNet9/CIFAR10.
@@
Additional Benchmarks. We give the detailed results of our evaluation in §A.3, including additional accuracy comparisons of WW-FL with the standard FL in all three modes of corruption, and the accuracy results for our Trimmed Mean variant.
6. Conclusion
@@
In this work, we presented WW-FL, a novel unified abstraction and framework for large-scale (hierarchical) federated learning that provides global model privacy, faster convergence, smaller attack surface, and better resilience against poisoning attacks than regular FL. We envision that WW-FL will lead to a paradigm shift in global FL system design and deployment. Future efforts can extend our foundation work with efficient instantiations, stronger security, and comprehensive evaluations with larger datasets and setups.
Acknowledgements
@@
This project received funding from the ERC under the EU’s Horizon 2020 research and innovation program (grant agreement No. 850990 PSOTI). It was co-funded by the DFG within SFB 1119 CROSSING/236615297 and GRK 2050 Privacy & Trust/251805230.
Appendix A Additional Details for WW-FL
This section provides additional details for our WW-FL framework, including motivation, security, and benchmarks.
A.1. Motivation for WW-FL
We elaborate on the motivation behind our WW-FL framework, which is specifically designed to facilitate the global deployment of FL. At a high level, WW-FL aims to address the diversity in trust levels among entities across different geographical locations (OECD, 2022; Statista Research Department, 2022). According to a survey (Statista Research Department, 2022), the global average trust in government was reported at $51\%$ in 2022. Furthermore, the degree of trust varies among different entities, such as law enforcement agencies and judicial systems (OECD, 2022).
When it comes to private companies, some countries have implemented strict laws to ensure the ethical behavior of entities operating within their borders. Non-compliance can result in substantial fines and even imprisonment (Husain, 2023; Hill, 2023). Recent examples include Meta receiving a 1.2 billion EUR fine from regulators in Ireland (Husain, 2023) and Didi Global facing a 1.19 billion EUR penalty from the Cyberspace Administration of China (Hill, 2023). On the other hand, several companies including well-known TalkTalk and Target either lost their customers’ data in major data breaches or were accused of not keeping customer information safe. This lead to a lack of trust among their clients and thereby face a huge drop in its revenue as well as loosing their customer base (Erin, 2022; Hibberd, 2022).
Given this disparity in trust, consider a consortium comprising three companies interested in training a machine learning (ML) model for medical diagnostics. Numerous studies have demonstrated the effectiveness of ML models in diagnosing various diseases, such as lung cancer, tumor detection, and liver segmentation (Rauniyar et al., 2023). However, these studies emphasize the need for diverse patient data. Simultaneously, the sensitivity and privacy of the training data must be preserved.
To address this, the consortium may collaborate with various federal, and local government agencies to obtain relevant training data and potentially compensating data providers. However, since the model is trained on sensitive medical data, the consortium members do not want the trained model to be disclosed to anyone. Additionally, the consortium should have the ability to use the trained model for inference while preserving privacy. This approach not only enhances the system’s privacy by safeguarding against breaches that might affect individual companies but also safeguards the financial investments of these companies by preventing unauthorized use of their trained model by third parties.
At first glance, implementing a Privacy-Preserving Machine Learning (PPML) solution may appear simplistic. In PPML, data owners securely distribute their data among the three companies within the consortium, ensuring that no individual company can access the data in the clear. The consortium employs MPC techniques to privately train a model on the shared data. However, this approach has several drawbacks. One significant drawback is the substantial computational power required within the consortium to handle the training process, particularly when dealing with millions of data owners in a global deployment. Furthermore, this approach necessitates a high level of trust from all data owners, as they must share their data among the consortium’s member companies. This trust requirement can be a significant hurdle to adoption. Additionally, there may be situations where data owners are willing to share their data but cannot do so due to government regulations preventing data from leaving their jurisdiction. This restriction not only limits data availability but also renders this approach unsuitable for our specific use case.
The next approach we can consider is using a traditional FL with secure aggregation. While this method has the potential to address the scalability problem, it comes with a significant drawback. In this approach, the trained model becomes publicly accessible to the data owners, which does not align with our specific scenario. Although there are works like (Froelicher et al., 2021; Sav et al., 2021), which protect the privacy of the global model using homomorphic encryption techniques, they may not be well-suited for our cross-device setting with a large number of data owners. Additionally, these techniques necessitate interaction among the clients and require extra measures for handling dropouts, as indicated in Tab. 1.
In light of the limitations of current methods, we have made a slight adjustment to the concept of FL in WW-FL. Specifically, we have expanded the notion of a trust zone, allowing data to be moved from the data owner to this larger, still-trusted area. For instance, a user in our scenario can securely share their data among three entities: the police, the court, and a private company, all operating within the same legal jurisdiction as the data owner. These selected entities will collaboratively train the model using privacy-enhancing technologies like MPC and HE. Furthermore, this approach has the added benefit of supporting edge nodes with limited computational resources, allowing them to contribute to the system. Importantly, since data owners can be assured that their data remains within their designated trust zone, this will likely encourage more users to participate, leading to a larger pool of training data and, ultimately, a better-trained model with improved global privacy.
Finally, it is important to understand that the concept of a trust-zone can differ depending on the specific region or context. For example, in one country, it may be relatively straightforward to identify three entities that data owners trust not to collude with each other. In contrast, there may be situations where it is challenging to find three or four entities that can provide an absolute guarantee of non-collusion. However, in such cases, data owners may be willing to accept a weaker assurance that not all entities will collude. WW-FL can accommodate this diversity in trust by performing PPML training independently for each of these scenarios, using the appropriate MPC protocol, and then combining their results through secure aggregation using global servers.
A.2. Security of WW-FL
This section provides security details of our WW-FL architecture from §3.1.
See 3.1
Proof Sketch
In the FL-Train algorithm, the initial step involves generating secret shares of the global model and distributing them among the global servers ($\mathcal{G}$). This step ensures correctness and privacy through the Share algorithm, a part of the $\mathcal{F}_{\textsf{MPC}}$ functionality. Once the global model is available in this secret-shared format, the next step is to securely redistribute it among servers within each cluster.
For example, consider the case where there are only two semi-honest global servers among which the global model is shared upon. Then, for an MPC cluster consisting of three servers (3PC), designed to tolerate one malicious corruption, the global servers employ the Reshare protocol in $\mathcal{F}_{\textsf{MPC}}$. This protocol converts their secret shares into the semantics suitable for the 3PC malicious cluster. Here again, the security relies upon the guarantees of the $\mathcal{F}_{\textsf{MPC}}$ ideal functionality.
The next step in FL-Train is for the clients to securely share their data with the cluster servers in their respective MPC clusters. This secure sharing is achieved through the Share protocol, which guarantees that the cluster servers will possess consistent shares of the data. Even in the event of malicious corruption, Share ensures that a malicious client cannot introduce inconsistent shares among the cluster servers. Note that a malicious client can tamper with its data before sharing it with an MPC cluster. However, this behavior is allowed in the $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ functionality, as $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ guarantees the privacy and correctness of the aggregation operation of the secret-shared data by the clients. The underlying secure aggregation algorithms (Agg) handle and address such attacks, which fall outside the scope of MPC. Nonetheless, $\mathcal{F}_{\textsf{MPC}}$ ensures the secure evaluation of these aggregation algorithms.
After the data and the current global model are distributed in secret-shared form among the cluster servers, they execute the Train algorithm to securely acquire the secret-shares of the trained cluster model. Subsequently, these cluster models are shared secretly among the global servers using the Reshare protocol. It is important to highlight that the security of these two steps is ensured by the $\mathcal{F}_{\textsf{MPC}}$ functionality. In the event of malicious corruption within the cluster, MPC protocols capable of tolerating such corruption will be employed. This will effectively detect any cheating attempt by a malicious cluster server with a very high probability.
After the cluster models have been transformed into secret shares, the global servers ($\mathcal{G}$) use the Agg algorithm to securely aggregate them and obtain the updated global model in secret-shared form. Note that the Agg algorithm, as mentioned before, plays a crucial role in defending against poisoning attacks from malicious clients and is outside the scope of MPC. However, it ensures the privacy and accuracy of the aggregation process, resulting in the correct global model. Finally, the global servers uses Reshare protocol to re-share the global model back to the cluster servers in their respective sharing semantics.
In summary, the correctness of Share, Reshare, Train, and Agg algorithms relies on the $\mathcal{F}_{\textsf{MPC}}$ functionality when instantiated correctly for the desired level of security. Therefore, with a very high probability, a malicious server within a cluster or at the global level cannot introduce errors into the system without being detected. At a high level, the underlying MPC protocols employ techniques such as error correction or replication in honest-majority settings, while dishonest-majority scenarios utilize message authentication codes to detect such errors. $\square$
A.3. Additional Benchmarks
@@
A.3.1. Additional Results for Q1
Fig. 7 shows validation accuracy of FL and WW-FL for image classification tasks over the full 2000 training iterations.
Figure 7. Validation accuracy for FL and $WW\mbox{-}FL$ for 2000 epochs (top: LeNet/MNIST, bottom: ResNet9/CIFAR10)
A.3.2. Additional Results for Q4
@@
WW-FL outperforms regular FL in nearly all settings and has more than 5% higher validation accuracy after 2000 rounds. In both regular FL and WW-FL, all three aggregation schemes show robustness against data poisoning in the realistic settings with 0.01 and 0.1 poison rate and equally-distributed as well as focused attacks. This is not the case for a poison rate of 0.2. Especially FedAvg as a non-robust aggregation scheme struggles to converge. Even robust aggregations cannot fully counteract the attacks. The same is true for the cluster-focused attack distribution: Here, all aggregations have spikes in their validation accuracy; only FLTrust manages to train a model comparable to the other settings. This is because FLTrust is designed to be robust against model poisoning, and the cluster-focused setting is the closest to a model-poisoning attack. We stress that the cluster-focused scenario is less practical in reality. Consider worldwide training that is organized into country-level clusters: it is extremely unlikely that an adversary can corrupt even the majority of a country’s population. Corruption in a production-level cross-device setting is below a rate of 0.001 (Shejwalkar et al., 2022), making our study of extreme cases with a 0.2 corruption rate a theoretical experiment.
Figure 8. Validation accuracy for ResNet9/CIFAR10 training with FedAvg, FLTrust, and trimmed mean for 2000 iterations under RLF, SLF, DLF, and TLF attacks in the equally-distributed setting.
Figure 9. Validation accuracy for ResNet9/CIFAR10 training with FedAvg, FLTrust, and trimmed mean for 2000 iterations under RLF, SLF, DLF, and TLF attacks in the focused setting.
Figure 10. Validation accuracy for ResNet9/CIFAR10 training with FedAvg, FLTrust, and trimmed mean for 2000 iterations under RLF, SLF, DLF, and TLF attacks in the cluster-focused setting.
Figure 11. Validation accuracy for LeNet/MNIST training with FedAvg, FLTrust, and trimmed mean for 2000 iterations under RLF, SLF, DLF, and TLF attacks in the equally-distributed setting.
Figure 12. Validation accuracy for LeNet/MNIST training with FedAvg, FLTrust, and trimmed mean for 2000 iterations under RLF, SLF, DLF, and TLF attacks in the focused setting.
Figure 13. Validation accuracy for LeNet/MNIST training with FedAvg, FLTrust, and trimmed mean for 2000 iterations under RLF, SLF, DLF, and TLF attacks in the cluster-focused setting.
Figure 14. Validation accuracy for ResNet9/CIFAR10 training with trimmed mean and our trimmed mean variant for 2000 iterations under RLF, SLF, DLF, and TLF attacks in the equally-distributed setting.
Figure 15. Validation accuracy for ResNet9/CIFAR10 training with trimmed mean and our trimmed mean variant for 2000 iterations under RLF, SLF, DLF, and TLF attacks in the focused setting.
Figure 16. Validation accuracy for ResNet9/CIFAR10 training with trimmed mean and our trimmed mean variant for 2000 iterations under RLF, SLF, DLF, and TLF attacks in the cluster-focused setting.

### v4

@@ 5.3 CrypTen Implementation Details for Trimmed Mean
When computing the proposed trimmed mean variant, we are only interested in the indices of the outliers and therefore only perform the swap operations on a list of indices. This is done to minimize the compute operations and thereby reduce time and communication. After identifying which gradients were most often detected as outliers and then computing the set of benign indices, we finally compute the sum over those while preserving privacy. The procedure is shown in Lst. 2.
@@
3 counts = crypten.stack([indices for _ in range(num_bins)], dim=1).eq(indices_list).sum(dim=0)
@@
5
6def exclude_topk_indices(self, indices, k):
@@
10
11benign_indices = self.exclude_topk_indices(
@@
14
15all_indices = crypten.cryptensor(torch.arange(0, n).repeat(benign_indices.size(0), 1)).t()
16eq_mask = all_indices.eq(benign_indices)
17agg_sum = gradients.matmul(eq_mask).sum(dim=1)
18
19return agg_sum / (n - self.exclude_topk)
@@
Table 10: Robust aggregation parameters.
Parameter FL $W$-FL
@@
FLTrust [CFLG21] uses a root data set to calculate the server model. To be most effective, the server model must be similar to the benign models and therefore the root dataset must be representative for the whole dataset. Therefore, we sample (like for all clients) 200 data points for that dataset.
5.4 Evaluation
We now empirically evaluate the impact of data-poisoning attacks on WW-FL considering different (robust) aggregation schemes. For this, we implement all four attacks—RLF [XXE12], TLF [TTGL20], SLF [FCJG20, SHKR22], DLF [SHKR22]—in our framework and add CrypTen-based implementations of three robust aggregation schemes to our prototype implementation. Using the setup described in §4, we answer the following questions next:
  • Q5:
@@ 5.4 Evaluation
Q5 – Attack Impact
We consider three poison rates (0.01, 0.1, 0.2) and distributions of attackers: in the equally-distributed setting, we assume that on expectation each cluster has the same number of malicious clients; in the focused setting, we assume that malicious clients are concentrated on as few clusters as possible while there is still an honest majority in each cluster (a standard assumption in FL); finally, in the unrealistic cluster-focused setting, we see what happens if we lift the honest-majority assumption and concentrate all malicious clients in as few clusters as possible. In Fig. 5, we study the effectiveness of the most powerful DLF [SHKR22] attack on both regular FL and WW-FL when training ResNet9 on CIFAR10 in the equally distributed and focused setting. Additional experimental outcomes, including evaluation over the cluster-focused setting, are given in §A.
Figure 5: Validation accuracy for FL and WW-FL when training ResNet9 on CIFAR10 with FedAvg, FLTrust, and trimmed mean as aggregation schemes under DLF attack for three different poison rates (top: equally distributed, bottom: focused setting).
@@
Q6 – Robust Aggregation in MPC
[…]
Table 11: Communication (Comm. in MB) and run-time (Time in seconds) for various aggregation schemes, including our Trimmed Mean (TM) variant with sample sizes 10, 100, and 1000 in CrypTen.
@@
Q7 – Trimmed Mean Variant
In Fig. 6, we compare the effectiveness of our TM variant to the original TM [YCRB18] for three sample sizes (10, 100, and 1000). It turns out that our heuristic approach barely reduces the effectiveness, even with aggressive parameters. In fact, in the focused setting, the TM variant outperforms the original. This is because our variant completely excludes updates of (poisoned) outliers, whereas in regular trimmed mean, those poisoned updates might still be considered for some coordinates. Results for all other settings are presented in §A.
Figure 6: Accuracy of trimmed mean (TM) and our variant (with sample sizes 10, 100, and 1000) against focused DLF attacks on WW-FL at 0.2 poison rate for ResNet9/CIFAR10.
@@
We provide additional results in §A, including additional accuracy comparisons of WW-FL with standard FL in all three modes of corruption, and the validation accuracy results for our Trimmed Mean variant.
6 Conclusion
@@
In this work, we presented WW-FL, a novel unified abstraction and framework for large-scale (hierarchical) federated learning that provides global model privacy, faster convergence, smaller attack surface, and better resilience against poisoning attacks than regular FL. We envision that WW-FL will lead to a paradigm shift in FL system design and deployment. Future efforts can extend our foundational work with more efficient instantiations, stronger security, and comprehensive evaluations with larger datasets and setups.
Acknowledgments
@@
This project received funding from the European Research Council (ERC) under the European Union’s research and innovation programs Horizon 2020 (PSOTI/850990) and Horizon Europe (PRIVTOOLS/101124778). It was co-funded by the Deutsche Forschungsgemeinschaft (DFG) within SFB 1119 CROSSING/236615297.
Appendix A Additional Benchmarks
@@
A.1 Additional Results for Q4
@@
WW-FL outperforms regular FL in nearly all settings and has more than 5% higher validation accuracy after 2000 rounds. In both regular FL and WW-FL, all three aggregation schemes show robustness against data poisoning in the realistic settings with 0.01 and 0.1 poison rate and equally-distributed as well as focused attacks. This is not the case for a poison rate of 0.2. Especially FedAvg as a non-robust aggregation scheme struggles to converge. Even robust aggregations cannot fully counteract the attacks. The same is true for the cluster-focused attack distribution: Here, all aggregations have spikes in their validation accuracy; only FLTrust manages to train a model comparable to the other settings. This is because FLTrust is designed to be robust against model poisoning, and the cluster-focused setting is the closest to a model-poisoning attack. We stress that the cluster-focused scenario is less practical in reality. Consider worldwide training that is organized into country-level clusters: it is extremely unlikely that an adversary can corrupt even the majority of a country’s population. Corruption in a production-level cross-device setting is below a rate of 0.001 [SHKR22], making our study of extreme cases with a 0.2 corruption rate a theoretical experiment.
Figure 7: Validation accuracy for ResNet9/CIFAR10 training with FedAvg, FLTrust, and trimmed mean for 2000 iterations under RLF, SLF, DLF, and TLF attacks in the equally-distributed setting.
Figure 8: Validation accuracy for ResNet9/CIFAR10 training with FedAvg, FLTrust, and trimmed mean for 2000 iterations under RLF, SLF, DLF, and TLF attacks in the focused setting.
Figure 9: Validation accuracy for ResNet9/CIFAR10 training with FedAvg, FLTrust, and trimmed mean for 2000 iterations under RLF, SLF, DLF, and TLF attacks in the cluster-focused setting.
Figure 10: Validation accuracy for LeNet/MNIST training with FedAvg, FLTrust, and trimmed mean for 2000 iterations under RLF, SLF, DLF, and TLF attacks in the equally-distributed setting.
Figure 11: Validation accuracy for LeNet/MNIST training with FedAvg, FLTrust, and trimmed mean for 2000 iterations under RLF, SLF, DLF, and TLF attacks in the focused setting.
Figure 12: Validation accuracy for LeNet/MNIST training with FedAvg, FLTrust, and trimmed mean for 2000 iterations under RLF, SLF, DLF, and TLF attacks in the cluster-focused setting.
Figure 13: Validation accuracy for ResNet9/CIFAR10 training with trimmed mean and our trimmed mean variant for 2000 iterations under RLF, SLF, DLF, and TLF attacks in the equally-distributed setting.
Figure 14: Validation accuracy for ResNet9/CIFAR10 training with trimmed mean and our trimmed mean variant for 2000 iterations under RLF, SLF, DLF, and TLF attacks in the focused setting.
Figure 15: Validation accuracy for ResNet9/CIFAR10 training with trimmed mean and our trimmed mean variant for 2000 iterations under RLF, SLF, DLF, and TLF attacks in the cluster-focused setting.